# Seg-Zero: Reasoning-Chain Guided Segmentation via Cognitive Reinforcement

### v1

@@ Abstract
Traditional methods for reasoning segmentation rely on supervised fine-tuning with categorical labels and simple descriptions, limiting its out-of-domain generalization and lacking explicit reasoning processes. To address these limitations, we propose Seg-Zero, a novel framework that demonstrates remarkable generalizability and derives explicit chain-of-thought reasoning through cognitive reinforcement. Seg-Zero introduces a decoupled architecture consisting of a reasoning model and a segmentation model. The reasoning model interprets user intentions, generates explicit reasoning chains, and produces positional prompts, which are subsequently used by the segmentation model to generate precious pixel-level masks. We design a sophisticated reward mechanism that integrates both format and accuracy rewards to effectively guide optimization directions. Trained exclusively via reinforcement learning with GRPO and without explicit reasoning data, Seg-Zero achieves robust zero-shot generalization and exhibits emergent test-time reasoning capabilities. Experiments show that Seg-Zero-7B achieves a zero-shot performance of 57.5 on the ReasonSeg benchmark, surpassing the prior LISA-7B by 18%. This significant improvement highlights Seg-Zero’s ability to generalize across domains while presenting an explicit reasoning process. Code is available at https://github.com/dvlab-research/Seg-Zero.
@@ 2.1 Reasoning in Large Models
In recent years, Large Language Models (LLMs) have exhibited remarkable reasoning capabilities. By extending the length of the Chain-of-Thought (CoT) reasoning process, OpenAI-o1 [27] introduces inference-time scaling, significantly enhancing its reasoning performance. In the research community, several studies have attempted to achieve test-time scaling through various approaches, including process-based reward models [39, 38, 20], reinforcement learning (RL) [15, 34], and search algorithms [10, 37]. Notably, the recent DeepSeek-R1 [11], which utilizes the GRPO [34] algorithm, achieves superior performance with only a few thousand RL training steps. Building on advancements in the LLMs community, several recent works have attempted to leverage the reasoning capabilities of MLLMs [16, 36]. For example, Open-R1-Multimodal [16] emphasizes mathematical reasoning, while R1-V [36] demonstrates exceptional performance in counting tasks. However, these works primarily address high-level reasoning and do not consider fine-grained pixel-level understanding of images. To fill this gap, our Seg-Zero is designed to enhance pixel-level reasoning through reinforcement learning.
@@ 2.2 Semantic Segmentation with Reasoning
Semantic segmentation aims to predict segmentation masks for specific classes. Numerous studies [33, 25, 21, 4, 5, 44, 1, 8], including DeepLab [6], MaskFormer [9] and SAM [14] have made significant progress in this task, making it a well-addressed problem. Instead of segmenting objects with explicit class labels, referring expression segmentation [13, 43] focuses on segmenting target objects based on short, explicit text queries. This task is more challenging as images often contain multiple instances of the same object class with varying attributes, requiring the model to identify and segment the instance that most closely matches the text query. LISA [17] advances this field further by introducing the reasoning segmentation task. In this task, the text queries are either more intricate or significantly longer, demanding models with strong reasoning capabilities to accurately interpret and segment the target objects.
@@
2.3 MLLMs for Segmentation
Since LISA [17, 41] introduced the ‘ <SEG >’ token to bridge the gap between MLLMs and segmentation models, several subsequent works [7, 32, 3] have explored the use of MLLMs for segmentation tasks. Most of these approaches, including OneTokenSegAll [3] and PixelLM [32], follow LISA’s paradigm by utilizing special tokens to connect MLLMs with segmentation models. However, this design necessitates extensive data to fine-tune both the MLLM and the segmentation decoder, and may even compromise the pixel precious of the original segmentation models. Our proposed Seg-Zero also employs a decoupled design for ease of adoption, while further leveraging the reasoning ability of MLLMs to achieve superior results.
@@ 3 Method
In this section, we introduce our Seg-Zero model and the associated reinforcement learning framework. We first describe how we address the segmentation problem in Section 3.1. Next, we present the architecture of the Seg-Zero in Section 3.2. Finally, we describe the reward functions (Section 3.3) and training details (Section 3.4) in the reinforcement learning framework.
@@ 3.1 Pipeline Formulation
Given an image $\mathbf{I}$ and a label $\mathbf{T}$ , the segmentation task aims to produce a binary segmentation mask $\mathbf{M}$ that accurately identifies the region corresponding to $\mathbf{T}$. The label $\mathbf{T}$ can vary in complexity, ranging from a simple class label (e.g., ”bird”), to a straightforward phrase (e.g., ”woman in blue”), or even to long and intricate expressions (e.g., ”The unusual thing in the image”). The latter two types of expressions require the model to perform reasoning to accurately segment the most relevant objects.
Inspired by recent advancements in the reasoning capabilities of large models [34, 36, 11], we leverage this ability to develop a pipeline for reasoning-based segmentation. Specifically, we decouple the reasoning process and the segmentation process. We first employ reinforcement learning to a MLLM to activate its reasoning ability, enabling it to generate reasoning process and produce accurate bounding box $\mathbf{B}$ and two points $\mathbf{P_{1},P_{2}}$ that best localize the target object. These bounding box and points are then used as prompts for SOTA segmentation models [14, 30] to produce fine-grained segmentation masks. Seg-Zero is trained using reinforcement learning, as illustrated in Figure 2.
@@ 3.2 Seg-Zero Model
Current MLLMs [40, 18, 45, 2, 24] exhibit impressive performance in processing multi-modal inputs but are unable to generate fine-grained segmentation masks. Conversely, modern segmentation models [14, 30] provides fine-grained segmentation ability but lack robust reasoning capabilities. To bridge this gap, we propose Seg-Zero, a framework that includes a reasoning model and a segmentation model. Additionally, we introduce the novel strategy to effectively activate the reasoning ability of MLLM within the framework. Its whole architecture is shown in Figure 3.
Reasoning Model. We employ Qwen2.5-VL [2] as our reasoning model $\mathcal{F}_{reason}$. While Qwen2.5-VL demonstrates exceptional performance in object detection by predicting the bbox, this region-level bbox is insufficient for providing more fine-grained pixel-level localization. Unlike object detection, segmentation requires a more precise understanding of pixel-level details, as multiple objects may exist within a single bounding box. Therefore, in addition to bounding box, we also incorporate points that lie within the target object to improve localization accuracy. During the reinforcement learning stage, the format rewards are employed to ensure the model generates structured outputs, which are subsequently processed by a post-processing function $\mathcal{G}$ to extract the bounding box $\mathbf{B}$ and the two points $\mathbf{P_{1},P_{2}}$ This process can be formulated as:
@@
Segmentation Model. Modern segmentation models [14, 30] accept various types of prompts, including bounding boxes and points, to generate accurate segmentation masks. We employ SAM2 [30] as our segmentation model $\mathcal{F}_{seg}$ due to its superior performance and efficient inference speed. Leveraging the bounding boxes and points provided by the reasoning model, the segmentation model can generate precise, fine-grained mask for the target object. This process can be formally expressed as:
@@
Test-time Reasoning. Reasoning is the crucial part in reasoning segmentation tasks. Inspired by DeepSeek-R1-Zero, we intentionally avoid using any explicit Chain-of-Thought (CoT) data to teach Seg-Zero reasoning skills. Instead, we aim to activate its reasoning capabilities from zero, enabling the model to autonomously generate a logical CoT before producing the final answer. To achieve this, we design a structured user prompt and a sophisticated reward mechanism that guides the reasoning model to follow specific instructions. As shown in Figure 4, this user prompt instructs Seg-Zero to analyze and compare objects in the image, beginning by generating a reasoning process, followed by the final answer in a predefined format.
@@ 3.3 Reward Functions
Segmentation Format Reward. Different from counting or other QA tasks, segmentation task is highly dependent on the format of the answer. We provide two types of segmentation format rewards: soft and strict. Under soft constrains, if the keywords bbox and points appear in the answer, and their corresponding values consist of four and two coordinates, respectively, the format is considered correct. Under strict constraints, the format is only considered correct if the model outputs exact keywords (e.g., bbox, points_1, points_2) in the required structure.
Bbox IoU Reward. This reward evaluates the IoU between the predicted bbox against the ground-truth bbox. A reward of 1 is assigned if their IoU greater than 0.5; otherwise, the reward is 0.
@@
Point L1 Reward. This reward evaluates the L1 distance between predicted points and the ground-truth points. We first determine whether the predicted points are inside the bounding box. Then the reward is set to 1 if the minimal distance between the predicted points and the ground-truth points is less than 100 pixels; otherwise, the reward is 0.
@@ 3.4 Training
We build the training data from public available segmentation datasets and train our Seg-Zero using GRPO algorithm.
Data Preparation. The training data is generated using the original mask annotations from existing referring expression segmentation datasets (e.g., RefCOCOg [43]). Based on the mask, we extract the leftmost, topmost, rightmost, and bottommost pixels of the mask to generate the bounding box $\mathbf{B}$. Additionally, we compute the center points of the two largest inscribe circles within the mask, denoted as $\mathbf{P_{1}}$ and $\mathbf{P_{2}}$ Consequently, the ground truth data comprises the bbox coordinates [$\mathbf{B}_{x1}$, $\mathbf{B}_{y1}$, $\mathbf{B}_{x2}$, $\mathbf{B}_{y2}$] and the coordinates of the two center points [$\mathbf{P_{1}}_{x}$, $\mathbf{P_{1}}_{y}$] and [$\mathbf{P_{2}}_{x}$, $\mathbf{P_{2}}_{y}$]. We do not incorporate any CoT processing into the training data. To ensure consistency across different datasets, all images are rescaled to a uniform resolution of 840x840 pixels.
GRPO. We do not include any reasoning data for a cold-start training process to teach the model’s reasoning ability. Instead, we let our Seg-Zero learn from zero. Specifically, we initiate training directly from the pre-trained Qwen2.5-VL-3B model, utilizing the aforementioned rewards and applying the GRPO algorithm [34]. We illustrate oue RL training process in Figure 2.
@@ 4 Experiment
4.1 Experimental Settings
Datasets. We training our Seg-Zero with only 9,000 samples adopted from RefCOCOg, using data preparation strategy mentioned in Section 3.4. The test data includes RefCOCO(+/g) [43] and ReasonSeg [17].
Implementation Details. We employ Qwen2.5-VL-3B [2] and SAM2-Large [30] as our default reasoning model and segmentation model, respectively. Seg-Zero is trained on an 8xH200 GPU server using the DeepSpeed [29] library. During training, we use a total batch size of 16 with a sampling number of 8 per training step. The initial learning rate is set to 1e-6 and the weight decay is 0.01.
4.2 SFT vs. RL
We compare the performance of SFT and RL. The baseline model is Qwen2.5-VL-3B + SAM2. For the non-CoT setting, we eliminate the thinking format reward, thus the model does not generate a CoT reasoning process before outputting the final answer. Our comparison includes both in-domain and OOD segmentation tasks [35, 26], as well as general QA tasks. The corresponding results are shown in Table 1, Figure 1 and Figure 5.
SFT vs. RL without CoT. From the first two rows in Table 1, we observe that on the in-domain dataset RefCOCOg, SFT achieves nearly the same performance as the baseline model. This may be because of the strong baseline performance of the original Qwen2.5-VL-3B. However, its performance significantly declines on the OOD ReasonSeg dataset, suggesting that SFT negatively impacts the model’s generalization ability. In contrast, comparing the first and third rows, we find that RL consistently improves performance on both in-domain and OOD datasets, demonstrating the effectiveness of RL. Besides, from Figure 5, we observe that the SFT model suffers from catastrophic forgetting of its original visual QA ability, while the RL model effectively preserves this capability.
RL without CoT vs. RL with CoT. From the last two rows in Table 1, we find that both RL and RL with CoT achieves superior performance on both the in-domain RefCOCOg and OOD ReasonSeg datasets, significantly outperforming the baseline. This indicates that RL effectively boost the models’ capabilies. However, with CoT, our Seg-Zero demonstrates even better performance compared to its counterparts without CoT, indicating the reasoning process enhances the models ability to handle OOD data samples. From Figure 5, it is noteworthy that the introduction of CoT reasoning leads to a slight performance improvement in visual QA tasks for models trained with RL.
@@ 4.3 Ablation Study
Soft vs. Hard Accurary Rewards. In Section 3.3, we describe the bbox IoU reward, the bbox L1 reward, and the point L1 reward. We apply specific thresholds to convert these metrics into binary rewards. Additionally, we conduct ablation studies on soft counterparts. For the bbox IoU reward, we directly use the IoU value as the soft reward. For the L1-based rewards, we define the soft reward as $1-\frac{L1\_dist}{max\{image\_size\}}$. From Table 3, we observe that while the soft reward achieves a minor improvement in ReasonSeg, it significantly underperforms compared to the hard reward on RefCOCOg.
@@
Soft vs. Strict Format Rewards. In Section 3.3, we introduce two types of segmentation format reward: the soft and strict. From Table 4, we find that the strict format reward significantly improves performance gain on OOD data in ReasonSeg. Through qualitative analysis of the training steps, we find that the strict format reward progresses slowly in the initial stages, as it is more challenging to sample formats that precisely match the strict criteria. However, as training step increases, model with strict format reward tend to output longer response.
@@
Changes in Completion Length. Figure 7 illustrates the trends in completion lengths across different model sizes. The results indicate that larger model tends to generate longer responses. As the training progresses, the minimal completion length gradually increases. However, there is a drop in average completion length during the initial few steps. By analyzing the output during the training process, we find that this occurs because the model initially prioritizes learning the correct output format, which often results in shorter responses. Once the format reward saturates, the model shifts its focus toward generating answers with higher accuracy, leading to longer and more detailed responses.
@@
Table 7: Referring expression segmentation results. We compare the cIoU in this table.

### v2

@@ Abstract
Traditional methods for reasoning segmentation rely on supervised fine-tuning with categorical labels and simple descriptions, limiting its out-of-domain generalization and lacking explicit reasoning processes. To address these limitations, we propose Seg-Zero, a novel framework that demonstrates remarkable generalizability and derives explicit chain-of-thought reasoning through cognitive reinforcement. Seg-Zero introduces a decoupled architecture consisting of a reasoning model and a segmentation model. The reasoning model interprets user intentions, generates explicit reasoning chains, and produces positional prompts, which are subsequently used by the segmentation model to generate precious pixel-level masks. We design a sophisticated reward mechanism that integrates both format and accuracy rewards to effectively guide optimization directions. Trained exclusively via reinforcement learning with GRPO and without explicit reasoning data, Seg-Zero achieves robust zero-shot generalization and exhibits emergent test-time reasoning capabilities. Experiments show that Seg-Zero-7B achieves a zero-shot performance of 57.5 on the ReasonSeg benchmark, surpassing the prior LISA-7B by 18%. This significant improvement highlights Seg-Zero’s ability to generalize across domains while presenting an explicit reasoning process. All code will be made publicly available for future research.
@@ 2.1 Reasoning in Large Models
In recent years, Large Language Models (LLMs) have exhibited remarkable reasoning capabilities. By extending the length of the Chain-of-Thought (CoT) reasoning process, OpenAI-o1 [27] introduces inference-time scaling, significantly improving its reasoning performance. In the research community, several studies have attempted to achieve test-time scaling through various approaches, including process-based reward models [39, 38, 20], reinforcement learning (RL) [15, 34], and search algorithms [10, 37]. In particular, the recent DeepSeek-R1 [11], which uses the GRPO [34] algorithm, achieves superior performance with only a few thousand RL training steps. Building on advances in the LLMs community, several recent works have attempted to leverage the reasoning capabilities of MLLMs [16, 36]. For example, Open-R1-Multimodal [16] emphasizes mathematical reasoning, while R1-V [36] shows exceptional performance in counting tasks. However, these works primarily address high-level reasoning and do not consider fine-grained pixel-level understanding of images. To fill this gap, our Seg-Zero is designed to enhance pixel-level reasoning through reinforcement learning.
@@ 2.2 Semantic Segmentation with Reasoning
Semantic segmentation aims at predicting segmentation masks for specific classes. Numerous studies [33, 25, 21, 4, 5, 44, 1, 8], including DeepLab [6], MaskFormer [9] and SAM [14] have made significant progress in this task, making it a well-addressed problem. Instead of segmenting objects with explicit class labels, referring expression segmentation [13, 43] focuses on segmenting target objects based on short, explicit text queries. LISA [17] advances this field further by introducing the reasoning segmentation task. In this task, text queries are either more intricate or longer, demanding models with strong reasoning capabilities to accurately interpret and segment the target objects.
2.3 MLLMs for Segmentation
Since LISA [17, 41] introduced the ‘ <SEG >’ token to bridge the gap between MLLMs and segmentation models, several subsequent works [7, 32, 3] have explored the use of MLLMs for segmentation tasks. Most of these approaches, including OneTokenSegAll [3] and PixelLM [32], follow LISA’s paradigm by using special tokens to connect MLLMs with segmentation models. However, this design necessitates extensive data to fine-tune both the MLLM and the segmentation decoder, and may even compromise the pixel precious of the original segmentation models. Our proposed Seg-Zero also employs a decoupled design for ease of adoption, while further leveraging the reasoning ability of MLLMs to achieve superior results.
@@ 3 Method
In this section, we introduce our Seg-Zero model and the associated reinforcement learning framework. We first describe how we address the segmentation problem in Section 3.1. Next, we present the architecture of the Seg-Zero in Section 3.2. Finally, we describe the reward functions (Section 3.3) and the training details (Section 3.4) in the reinforcement learning framework.
@@ 3.1 Pipeline Formulation
Given an image $\mathbf{I}$ and a label $\mathbf{T}$ , the segmentation task aims to produce a binary segmentation mask $\mathbf{M}$ that accurately identifies the region corresponding to $\mathbf{T}$. The label $\mathbf{T}$ can vary in complexity, ranging from a simple class label (e.g., “bird”), to a straightforward phrase (e.g., “woman in blue”), or even to long and intricate expressions (e.g., “The unusual thing in the image”). The latter two types of expression require the model to perform reasoning to accurately segment the most relevant objects.
Inspired by recent advancements in the reasoning capabilities of large models [34, 36, 11], we leverage this ability to develop a pipeline for reasoning-based segmentation. Specifically, we decouple the reasoning process and the segmentation process. We first employ reinforcement learning to an MLLM to activate its reasoning ability, enabling it to generate the reasoning process and produce accurate bounding box $\mathbf{B}$ and two points $\mathbf{P_{1},P_{2}}$ that best localize the target object. These bounding box and points are then used as prompts for SOTA segmentation models [14, 30] to produce fine-grained segmentation masks. Seg-Zero is trained using reinforcement learning, as illustrated in Figure 2.
@@ 3.2 Seg-Zero Model
Current MLLMs [40, 18, 45, 2, 24] exhibit impressive performance in processing multi-modal inputs but are unable to generate fine-grained segmentation masks. Conversely, modern segmentation models [14, 30] provide fine-grained segmentation ability but lack robust reasoning capabilities. To bridge this gap, we propose Seg-Zero, a framework that includes a reasoning model and a segmentation model. Additionally, we introduce the novel strategy to effectively activate the reasoning ability of MLLM within the framework. Its whole architecture is shown in Figure 3.
Reasoning Model. We employ Qwen2.5-VL [2] as our reasoning model $\mathcal{F}_{reason}$. Although Qwen2.5-VL demonstrates exceptional performance in object detection by predicting the bbox, this region-level bbox is insufficient to provide more fine-grained pixel-level localization. Unlike object detection, segmentation requires a more precise understanding of pixel-level details, as multiple objects may exist within a single bounding box. Therefore, in addition to the bounding box, we also incorporate points that lie within the target object to improve localization accuracy. During the reinforcement learning stage, the format rewards are employed to ensure the model generates structured outputs, which are subsequently processed by a post-processing function $\mathcal{G}$ to extract the bounding box $\mathbf{B}$ and the two points $\mathbf{P_{1},P_{2}}$ This process can be formulated as follows:
@@
Segmentation Model. Modern segmentation models [14, 30] accept various types of prompt, including bounding boxes and points, to generate accurate segmentation masks. We employ SAM2 [30] as our segmentation model $\mathcal{F}_{seg}$ due to its superior performance and efficient inference speed. Leveraging the bounding boxes and points provided by the reasoning model, the segmentation model can generate a precise, fine-grained mask for the target object. This process can be formally expressed as follows:
@@
Test-time Reasoning. Reasoning is the crucial part in reasoning segmentation tasks. Inspired by DeepSeek-R1-Zero, we intentionally avoid using any explicit Chain-of-Thought (CoT) data to teach Seg-Zero reasoning skills. Instead, we aim to activate its reasoning capabilities from zero, enabling the model to autonomously generate a logical CoT before producing the final answer. To achieve this, we design a structured user prompt and a sophisticated reward mechanism to guide the reasoning model toward the correct optimization direction. As shown in Figure 4, the user prompt instructs Seg-Zero to analyze and compare objects in the image, beginning by generating a reasoning process, followed by the final answer in a pre-defined format. The reward mechanism then evaluates the answers and directs the optimization process, as illustrated in Figure 2.
@@ 3.3 Reward Functions
Segmentation Format Reward. Different from counting or other QA tasks, the segmentation task is highly dependent on the format of the answer. We provide two types of segmentation format rewards: soft and strict. Under soft constraints, if the keywords bbox and points appear in the answer, and their corresponding values consist of four and two coordinates, respectively, the format is considered correct. Under strict constraints, the format is only considered correct if the model outputs exact keywords (e.g., bbox, points_1, points_2) in the required structure.
Bbox IoU Reward. This reward evaluates the IoU between the predicted bbox and the ground-truth bbox. A reward of 1 is assigned if their IoU greater than 0.5; otherwise, the reward is 0.
@@
Point L1 Reward. This reward evaluates the L1 distance between the predicted points and the ground-truth points. We first determine whether the predicted points are inside the bounding box. Then the reward is set to 1 if the minimal distance between the predicted points and the ground-truth points is less than 100 pixels; otherwise, the reward is 0.
@@ 3.4 Training
We build the training data from publicly available segmentation datasets and train our Seg-Zero using the GRPO algorithm.
Data Preparation. The training data is generated using the original mask annotations from existing referring expression segmentation datasets (e.g., RefCOCOg [43]). Based on the mask, we extract the leftmost, topmost, rightmost, and bottommost pixels of the mask to generate the bounding box $\mathbf{B}$. Additionally, we compute the center points of the two largest inscribe circles within the mask, denoted as $\mathbf{P_{1}}$ and $\mathbf{P_{2}}$ Consequently, the ground truth data comprises the bbox coordinates [$\mathbf{B}_{x1}$, $\mathbf{B}_{y1}$, $\mathbf{B}_{x2}$, $\mathbf{B}_{y2}$] and the coordinates of the two center points [$\mathbf{P_{1}}_{x}$, $\mathbf{P_{1}}_{y}$] and [$\mathbf{P_{2}}_{x}$, $\mathbf{P_{2}}_{y}$]. We do not incorporate any CoT processing into the training data. To ensure consistency, all images are rescaled to a uniform resolution of 840x840 pixels.
GRPO. We do not include any reasoning data for a cold-start training process to teach the model’s reasoning ability. Instead, we let our Seg-Zero evolve from zero. Specifically, we initiate training directly from the pre-trained Qwen2.5-VL-3B model, utilizing the aforementioned rewards and applying the GRPO algorithm [34]. We illustrate our RL training process in Figure 2.
@@ 4 Experiment
4.1 Experimental Settings
Datasets. We training our Seg-Zero with only 9,000 samples adopted from RefCOCOg, using the data preparation strategy mentioned in Section 3.4. The test data includes ReasonSeg [17] and RefCOCO(+/g) [43].
Implementation Details. We employ Qwen2.5-VL-3B [2] and SAM2-Large [30] as our default reasoning model and segmentation model, respectively. Seg-Zero is trained using the DeepSpeed [29] library. During training, we use a total batch size of 16 with a sampling number of 8 per training step. The initial learning rate is set to 1e-6 and the weight decay is 0.01.
4.2 SFT vs. RL
We compare the performance of SFT and RL. The baseline model is Qwen2.5-VL-3B + SAM2-Large. For the non-CoT setting, we eliminate the thinking format reward, thus the model does not generate a CoT reasoning process before outputting the final answer. Our comparison includes both in-domain and OOD segmentation tasks [35, 26], as well as general QA tasks. The corresponding results are shown in Table 1, Figure 1 and Figure 5.
@@ 4.1 Experimental Settings
SFT vs. RL without CoT. From the first two rows in Table 1, we observe that on the in-domain dataset RefCOCOg, SFT achieves nearly the same performance as the baseline model. This may be due to the strong baseline performance of the original Qwen2.5-VL-3B. However, its performance significantly declines on the OOD ReasonSeg dataset, suggesting that SFT negatively impacts the model’s generalization ability. In contrast, comparing the first and third rows, we find that RL consistently improves performance on both in-domain and OOD datasets, demonstrating the effectiveness of RL. Besides, from Figure 5, we observe that the SFT model suffers from catastrophic forgetting of its original visual QA ability, while the RL model effectively preserves this capability.
RL without CoT vs. RL with CoT. From the last two rows in Table 1, we find that both RL and RL with CoT achieve superior performance on both the in-domain RefCOCOg and OOD ReasonSeg datasets, significantly outperforming the baseline. This indicates that RL effectively boosts the models’ capabilities. However, with CoT, our Seg-Zero demonstrates even better performance compared to its counterparts without CoT, indicating that the reasoning process enhances the model’s ability to handle OOD data samples. From Figure 5, it is noteworthy that the introduction of CoT reasoning leads to a slight performance improvement in visual QA tasks for models trained without CoT.
@@ 4.3 Ablation Study
Model coef RefCOCOg ReasonSeg
Seg-Zero 1e-3 73.7 52.7
Seg-Zero 5e-3 73.6 53.8
Seg-Zero 1e-2 70.6 53.3
Seg-Zero 5e-2 66.6 50.8
Table 3: Ablation on the KL loss coefficient. This coefficient balance ‘pre-existing knowledge’ and ‘new knowledge’. Higher coefficient causes performance degradation.
KL Loss Coefficient. The KL loss coefficient balances the model’s ‘pre-existing knowledge’ with ‘new knowledge’. Table 3 presents the performance variations across different KL loss coefficients. We find that a coefficient of 5e-3 performs optimally on both in-domain and OOD data. A higher coefficient leads to performance degradation.
Model Number RefCOCOg ReasonSeg
Seg-Zero 4 66.8 52.0
Seg-Zero 8 73.6 53.8
Seg-Zero 16 74.1 54.7
Table 4: Ablation on number of samples. A larger sample number leads to better performance.
Number of Samples. We investigate the impact of the number of samples during the sampling stage. As shown in Table 4, we observe that as the number of samples increases, the model achieves better performance on both in-domain and out-of-distribution (OOD) data. This is reasonable because a larger number of samples expands the exploration space, enabling the model to identify more effective optimization directions.
Model Example RefCOCOg ReasonSeg
Seg-Zero × 72.1 49.4
Seg-Zero 73.6 53.8
Table 5: Ablation on the the user prompt. Including an example in the user prompt is crucial.
User Prompt Sensitivity. The last two rows of Figure 4 show that we include output examples in the user prompt. We investigate the impact of this example in Table 5 and observe that its inclusion significantly enhances the model’s performance. Through analysis of the output, we find that models without this example often fail to generate a reasoning process in their responses.
@@
Soft vs. Hard Accuracy Rewards. In Section 3.3, we describe the bbox IoU reward, the bbox L1 reward, and the point L1 reward. We apply specific thresholds to convert these metrics into binary rewards. Additionally, we conduct ablation studies on soft counterparts. For the bbox IoU reward, we directly use the IoU value as the soft reward. For L1-based rewards, we define the soft reward as 1L1 $1-\dfrac{\text{L1 dist}}{\max\{\text{image size}\}}$. From Table 6, we observe that while the soft reward achieves a minor improvement on ReasonSeg, it significantly underperforms compared to the hard reward on RefCOCOg.
@@
Soft vs. Strict Format Rewards. In Section 3.3, we introduce two types of segmentation format rewards: the soft and strict. From Table 7, we find that the strict format reward significantly improves performance gain on OOD data in ReasonSeg. Through qualitative analysis of the training steps, we find that the strict format reward progresses slowly in the initial stages, as it is more challenging to sample formats that precisely match the strict criteria. However, as training step increases, model with strict format reward tend to output longer response.
@@
Changes in Completion Length. Figure 6 illustrates the trends in completion lengths between different model sizes. The results indicate that a larger model tends to generate longer responses. As training progresses, the minimal completion length gradually increases. However, there is a drop in average completion length during the initial few steps. By analyzing the output during the training process, we find that this occurs because the model initially prioritizes learning the correct output format, which often results in shorter responses. Once the format reward saturates, the model shifts its focus to generating answer with higher accuracy, leading to longer and more detailed responses. Supplementary materials provide more analysis.
@@
Table 10: Referring expression segmentation results. We compare cIoU in this table.